# Is In-Context Learning Sufficient for Instruction Following in LLMs?

### v1

@@ Abstract
In-context learning (ICL) allows LLMs to learn from examples without changing their weights, which is a particularly promising capability for long-context LLMs that can potentially learn from many examples. Recently, Lin et al. (2024) proposed Urial, a method using only three in-context examples to align base LLMs, achieving non-trivial instruction following performance. In this work, we show that, while effective, ICL alignment with Urial still underperforms compared to instruction fine-tuning on established benchmarks such as MT-Bench and AlpacaEval 2.0 (LC), especially with more capable base LMs. Unlike for tasks such as classification, translation, or summarization, adding more ICL demonstrations for long-context LLMs does not systematically improve instruction following performance. To address this limitation, we derive a greedy selection approach for ICL examples that noticeably improves performance, yet without bridging the gap to instruction fine-tuning. Finally, we provide a series of ablation studies to better understand the reasons behind the remaining gap, and we show how some aspects of ICL depart from the existing knowledge and are specific to the instruction tuning setting. Overall, our work advances the understanding of ICL as an alignment technique. We provide our code at https://github.com/tml-epfl/icl-alignment.
Machine Learning, ICML
@@ 1 Introduction
The large-scale pre-training phase allows Large Language Models (LLMs) to acquire extensive knowledge and capabilities. However, these base models struggle to follow instructions directly from prompts, necessitating further fine-tuning to be used as conversational models. Traditional alignment methods include Supervised Fine-Tuning (SFT) on instruction datasets (IFT, Chung et al., 2022) or preference data (DPO and KTO, Rafailov et al., 2023; Ethayarajh et al., 2024), and reinforcement learning (RLHF, Ouyang et al., 2022). Inspired by Brown et al. (2020) who showed that LLMs can learn from demonstrations provided as part of the input—a concept known as in-context learning (ICL)—Lin et al. (2024) have recently studied the feasibility of in-context alignment. They found that including merely three carefully selected question-answer pairs in the prompt is sufficient to make base models follow instructions and interact with users at a similar level to instruction-finetuned models on their own benchmark. The possibility of customizing base models via ICL without changing weights is potentially game-changing due to its simplicity and flexibility.
In this work, we evaluate the Urial prompt strategy proposed by Lin et al. (2024) across several base models, including GPT-4-Base [^1] We received access to the base GPT-4 model via the OpenAI Researcher Access Program., on established instruction following benchmarks like MT-Bench (Zheng et al., 2023) and AlpacaEval 2.0 (Li et al., 2023). Although Urial achieves reasonable performance, it still lags behind instruction-finetuned models. In response, we test various strategies to improve Urial, particularly leveraging models with extensive context windows which allow for longer in-context prompts. First, we study the effect of many-shot in-context learning by adding high-quality demonstrations from existing instruction datasets. This approach, however, leads only to limited improvements compared to Urial, without a clear improvement trend for a large number of in-context examples. This is in contrast to many-shot ICL for tasks like summarization, translation, or classification, where providing many examples is clearly beneficial (Agarwal et al., 2024; Bertsch et al., 2024). Then we propose a simple greedy algorithm to select the in-context examples which optimize the MT-Bench score. This selection scheme allows us to nearly close the gap between Mistral-7b-v0.2 and Mistral-7b-Instruct-v0.2 by adding two additional demonstrations to the Urial prompt. Finally, we revisit the findings of Min et al. (2022) regarding what makes ICL effective: unlike for classification and multiple choice tasks, providing questions with correct answers is crucial for the effectiveness of ICL on instruction-following tasks. Furthermore, adding incorrect demonstrations further degrades the model performance. Overall, our work provides a more nuanced picture of ICL as an alignment technique compared to previous works.
2 Systematic evaluation of Urial
First, we compare the performance of several base models with the Urial in-context prompt to that of their instruction fine-tuned versions. Table 1 shows the results on MT-Bench (Zheng et al., 2023) which is one of the most popular benchmarks for instruction following ability of LLMs. We note, however, that Lin et al. (2024) originally tested Urial on their own dataset, in which only $8\%$ of the examples are obtained from MT-Bench. We compare several LLMs of different sizes and capabilities, ranging from Llama-2-7B (Touvron et al., 2023) to the base GPT-4 model (OpenAI, 2023). For simplicity, we consider the official instruction-tuned LLMs, i.e., provided from the same source of the base models, even if there exist several third-party fine-tuned versions which can achieve higher scores. We observe that base models with Urial achieve competitive performance on the 1st-turn score, but cannot match the performance of their instruction fine-tuned counterparts in all cases except for Llama-2-70B and Mistral-7B-v0.1 (note that both were originally included in Lin et al. (2024), unlike most others). Conversely, the 2nd-turn score with Urial is significantly worse than for instruction-tuned LLMs: we hypothesize that this is because no multi-turn demonstrations are given in context. Indeed, Zhou et al. (2023) show that, even in IFT, having a few multi-turn training examples significantly improve the performance of the model on multi-turn conversations. Because of this, in the rest of the paper, we focus on the 1st-turn score and single-round conversations, and leave extending ICL for multi-round conversations to future work. We detail the breakdown of the results from Table 1 over categories in Table 5.
Table 1: Systematic comparison of Urial to finetuned models on MT-Bench across different base LLMs. * denotes the result taken from the Urial GitHub repository.
@@ 2 Systematic evaluation of Urial
3 Many-shot ICL for instruction following
In the following, we explore whether we can close the gap between ICL and fine-tuning on instruction following tasks. In particular, we focus on Mistral-7B-v0.2 (Jiang et al., 2023) since (a) both Urial and the aligned model achieve competitive performance in Table 1, and (b) this base model has a 32k context window which allows us to use many ICL examples. Moreover, we measure performance with both the MT-Bench 1st-turn score and the AlpacaEval 2.0 win rate, and report the results in Table 2. More details can be found in App. A.
Figure 1: The results of scaling experiments with different sampling strategies on Mistral-7B-v0.2-32k. The black dashed line refers to the 1st-turn MT-Bench score of Mistral-7B-Instruct-v0.2. ICL with more demonstrations does not bridge the gap between the base model and its aligned counterpart. Surprisingly, with only three examples from Urial and questions without answers, the base model always achieves better instruction-following performance than with the full question-answer pairings.
Experimenting with different selection criteria. First, we test adding random demonstrations (after simple pre-filtering) to the three already included in Urial. For this, we score the quality of the examples in Evol-Instruct-70k (Xu et al., 2024) with GPT-3.5-Turbo, and randomly sample demonstrations among the highest quality instructions (those with score at least 4.5 on a 0-5 scale, see Fig. 4). Table 2 shows that adding 17 or 37 random examples (which makes 20 and 40 examples in total) provides only moderate improvements, around $+0.3$, and Fig. 1 indicates that such an increase quickly reaches a plateau (see the Evol-Instruct-Random curve), without clear benefits from scaling up to 50 demonstrations. We also test a more refined strategy: we sample instructions from OpenHermes (Teknium, 2023), which includes annotations on the topic of each example, to ensure diversity between categories. This approach is however worse than Urial, regardless of the number of ICL examples used (see the OpenHermes-Diversity curve in Fig. 1). Finally, inspired by the unsupervised ICL of Agarwal et al. (2024), we re-use the OpenHermes examples, this time without the answers (questions only) and adding instead the three Urial complete demonstrations (both questions and answers). This strategy significantly improves the MT-Bench score (from 7.0 to 7.5, see the OpenHermes-Unsupervised curve) and is slightly better than using the random examples from Evol-Instruct.
Figure 2: The distribution of the 1st-turn MT-Bench score (GPT-4-Turbo as judge) on Mistral-7B-v0.2-32k obtained by adding 100 instructions from Evol-Instruct-70k as a 4th demonstration to Urial. The black dashed line refers to the 1st-turn MT-Bench score of the base model with Urial. A majority of the 4th examples have positive contributions to the model’s instruction-following performance.
Adding examples based on greedy search. Next, we try to greedily maximize the MT-Bench score by testing 100 high-quality instructions from Evol-Instruct-70k as the 4th additional example to Urial. For each resulting ICL prompt, we compute the MT-Bench score with GPT-4-Turbo instead of GPT-4 as the former is faster, cheaper, and mitigates potential overfitting to the exact benchmark score. In most cases, there is an improvement, although it is often small (see the score distribution in Fig. 2). Computing the true MT-Bench (GPT-4 as judge) with the best performing point as the 4th ICL example yields a significant improvement over Urial, from 6.99 to 7.46 (Table 2). We then repeat this procedure sequentially to find a 5th and a 6th demonstration while restricting the search space at each round (by keeping only the instructions leading to a high enough MT-Bench score) to reduce the computational cost. As shown in Table 2, this strategy achieves 7.69 MT-Bench score, almost closing the gap with Mistral-7B-Instruct-v0.2 (8.06). In addition, it even improves AlpacaEval 2.0 win rate from 8.09 of Urial to 9.22 without directly optimizing for it. This improvement indicates that the found IC prompt does not completely overfit to the MT-Bench score. Finally, we use the scores obtained in the first round of the greedy search to scale-up the number of ICL examples: the Evol-Instruct-Greedy curve in Fig. 1 is obtained by progressively adding the instructions with the highest scores. However, adding demonstrations in this way beyond the 6 examples does not consistently improve the MT-Bench score.
Takeaways. These results show that, unlike in the setting of Agarwal et al. (2024); Bertsch et al. (2024), simply scaling the number of ICL examples is not sufficient to consistently improve the instruction following performance. However, a greedy search approach can help to bridge the gap to the results of fine-tuned models. We highlight that our greedy search is performed with a limited computational cost, i.e., the set of candidate examples is small and the judge is slightly suboptimal (GPT-4-Turbo instead of GPT-4). We expect that more effective ICL prompts could be found with additional resources.

### v2

@@ Abstract
In-context learning (ICL) allows LLMs to learn from examples without changing their weights: this is a particularly promising capability for long-context LLMs that can potentially learn from many examples. Recently, Lin et al. (2024) proposed Urial, a method using only three in-context examples to align base LLMs, achieving non-trivial instruction following performance. In this work, we show that, while effective, ICL alignment with Urial still underperforms compared to instruction fine-tuning on the established benchmark MT-Bench, especially with more capable base LLMs. We then uncover the most relevant elements for successful in-context alignment, finding the crucial role of the decoding parameters. Based on these insights, we show that the approach of Urial can indeed be improved by adding high-quality, possibly carefully selected via greedy search, demonstrations in context, getting closer to the performance of instruct models. Finally, we provide the first, to our knowledge, systematic comparison of ICL and instruction fine-tuning (IFT) for instruction following in the low data regime, where ICL can be a viable alternative to IFT. Overall, our work advances the understanding of ICL as an alignment technique and its relationship to IFT. We provide our code at https://github.com/tml-epfl/icl-alignment.
@@ 1 Introduction
The large-scale pre-training phase allows Large Language Models (LLMs) to acquire extensive knowledge and capabilities (Bubeck et al., 2023). However, these base models struggle to follow instructions directly from prompts, necessitating further fine-tuning to be used as conversational models. Traditional alignment methods include Supervised Fine-Tuning (SFT) on instruction datasets (IFT, Wei et al., 2021; Chung et al., 2022) or preference data (DPO, KTO, IPO, and ORPO, Rafailov et al., 2023; Ethayarajh et al., 2024; Azar et al., 2024; Hong et al., 2024), and reinforcement learning (RLHF and RLAIF, Ouyang et al., 2022; Bai et al., 2022). These approaches typically involve multiple rounds of refinement of the instructed model (Touvron et al., 2023), i.e. high computational cost, and need a large amount of annotated data like human preferences, which might be difficult to collect. However, a line of work (Taori et al., 2023; Chen et al., 2023; Zhou et al., 2023; Lee et al., 2023; Zhao et al., 2024; Kaur et al., 2024) has suggested that IFT on a small amount of high quality instruction-following examples, even only 1000, can be sufficient to match the performance of more complex alignment mechanisms. In particular, Zhou et al. (2023) introduced the Superficial Alignment Hypothesis, stating that LLMs acquire all their capabilities during pre-training, and fine-tuning only allows the models to better access such knowledge when interacting with users (Gudibande et al., 2023; Duan et al., 2023). Using such small instruction datasets for IFT has the clear advantage of significantly reducing the cost of model alignment.
@@
Inspired by Brown et al. (2020) who showed that LLMs can learn from demonstrations provided as part of the input—a concept known as in-context learning (ICL)—Lin et al. (2024) have recently studied the feasibility of in-context alignment (Han, 2023; Li et al., 2023b). They found that including merely three carefully selected question-answer pairs in the prompt is sufficient to make base models follow instructions and interact with users at a similar level to instruction-fine-tuned models on their own benchmark. This fact strongly validates the idea that alignment can be achieved at low computational cost and with few examples, as well the Superficial Alignment Hypothesis, as it does not even require to modify the model weights. Moreover, in-context alignment is especially appealing since it opens up the possibility of customizing base models via ICL, i.e., without any fine-tuning, which is potentially game-changing due to its simplicity and flexibility. While the results of Lin et al. (2024) are promising, they are restricted to a small number of base models and their own instruction following dataset.
Table 1: Systematic comparison of Urial to aligned models on MT-Bench across different base LLMs. For several recent model families, we compare the performance on instruction following tasks of the base LLMs plus Urial (i.e. in-context alignment) to that of the instruct models, fine-tuned with sophisticated techniques like supervised instruction fine-tuning and RLHF. In most cases, the fine-tuned models outperform Urial. * denotes the result taken from the Urial GitHub repository.
@@
Analysis of in-context alignment. In this work, we extend the evaluation of the Urial prompt strategy proposed by Lin et al. (2024) across several base models, including GPT-4-Base [^1] We received access to the base GPT-4 model via the OpenAI Researcher Access Program. and on established instruction following benchmark MT-Bench (Zheng et al., 2023), see Table 1. First, we show that, although Urial achieves reasonable performance, it still lags behind instruction-fine-tuned models. Next, to better understand the success but also weaknesses of in-context alignment, we analyze which are the most relevant ingredients in Urial. We find that the decoding parameters in the LLMs generation (temperature, sampling scheme, etc.) may crucially influence the performance, and the optimal configuration allows even base models to achieve good scores on MT-Bench. Moreover, we show that randomly sampled triplets of examples from the highly curated instruction dataset SkillMix (Kaur et al., 2024) can achieve similar, or even better, results than Urial when used for in-context alignment.
Many-shot in-context alignment. Next, we test various strategies to improve in-context alignment, leveraging recent models with extensive context windows which allow for longer in-context prompts. In particular, we study the effect of many-shot in-context learning by adding high-quality demonstrations from existing instruction datasets. Unlike what suggested by Lin et al. (2024), this approach can improve upon Urial, but only when using high-quality examples, as those from SkillMix. However, it is still not sufficient to fully close the gap with aligned LLMs, as the performance plateaus after 10-30 in-context examples. This behavior is in contrast to many-shot ICL for tasks like summarization (Narayan et al., 2018), translation (Costa-jussà et al., 2022), or classification (Li et al., 2024), where providing many examples is clearly beneficial (Agarwal et al., 2024; Bertsch et al., 2024). We further test a simple greedy algorithm to select the in-context examples which optimize the MT-Bench score. This selection scheme outperforms, with 1 to 3 additional demonstrations, the many-shot approach with random samples, and allows to further reduce the distance from in-context aligned models to fine-tuned models.
ICL vs IFT for alignment. While our experiments suggest that ICL cannot match the instruction-following performance of models aligned through costly fine-tuning, possibly using preference data, it remains an open question whether ICL can compete with or outperform IFT in the low-data regime. We provide an extensive evaluation on multiple LLMs and datasets of both approaches when varying the question-answer dataset size between 3 and 4000 examples. We observe that, with high-quality data, ICL and IFT achieve almost identical 1st-turn MT-Bench score. Surprisingly, in the 2nd-turn score, IFT clearly outperforms ICL, with ICL performing even worse than the base model. This suggests that ICL suffers from some heavy overfitting to the style of the examples shown in context.
Overall, these results, complemented by several ablation studies, provide a more nuanced picture of ICL as an alignment technique compared to previous works. Moreover, our comparison of ICL to IFT when using the same data bridges a gap in the literature about understanding these orthogonal approaches for adapting pre-trained LLMs into conversational models.
2 Uncovering the Limits of ICL for Instruction Following
In the following, we provide an in-depth analysis which aims at (1) systematically comparing the performance of base models plus Urial to that of aligned models (Sec. 2.1), (2) understanding which components of Urial (and in-context alignment in general) are most important for its effectiveness (Sec. 2.2), (3) testing the influence of question-answers format in our task (Sec. 2.3).
2.1 Systematic evaluation of Urial
Lin et al. (2024) propose to use a prompt with as few as three constant stylistic examples and a carefully designed set of rules to align base LLMs with ICL. The highly curated examples always begin with affirming the user query, then proceed to answer the query by enumerating necessary items and steps, and conclude with an engaging summary. The set of rules first elucidate the scenario and format of the subsequent interaction between the human user and the AI assistant, then outline multiple qualitative aspects, including helpfulness, honesty, and harmlessness, which the base LLMs should adhere to when answering queries.
We compare the performance of several base models with the Urial in-context prompt to that of their instruction fine-tuned versions. Table 1 shows the results on MT-Bench (Zheng et al., 2023), one of the most popular benchmarks for instruction following ability of LLMs. We note that Lin et al. (2024) originally tested Urial on their own dataset, in which only $8\%$ of the examples are obtained from MT-Bench. We compare several LLMs of different sizes and capabilities, ranging from Llama-2-7B (Touvron et al., 2023) to the base GPT-4 model (OpenAI, 2023). For simplicity, we consider the official instruction-tuned LLMs, i.e., provided from the same source of the base models, even if there exist third-party fine-tuned versions which can achieve higher scores. We observe that base models with Urial achieve competitive performance on the 1st-turn score, but cannot match that of their instruction fine-tuned counterparts in all cases except for Llama-2-70B and Mistral-7B-v0.1 (both originally included in Lin et al. (2024), unlike most others). Conversely, the 2nd-turn score with Urial is significantly worse than for instruction-tuned LLMs: we hypothesize that this is because no multi-turn demonstrations are given in context. Indeed, Zhou et al. (2023) show that, even in IFT, having a few multi-turn training examples significantly improve the performance of the model on multi-turn conversations. Because of this fact, in the rest of the paper, we mainly focus on the 1st-turn score and single-round conversations, and track of 2nd-turn performance for completeness. Finally, we detail the breakdown of the results over categories in Table 7 (see App. C.3 for details).
2.2 What matters for in-context alignment via Urial?
(a) Base model
(b) Base model with Urial
(c) Instruct model
Figure 1: Effect of decoding parameters on the 1st-turn MT-Bench scores across models. We vary temperature and top-$p$, with a fixed repetition penalty of 1.15 as used in the Urial codebase. The heatmaps show that the answering quality of the base model (Mistral-7B-v0.2) with and without Urial is sensitive to the decoding schemes. Conversely, the performance of the instruct model is robust to varying the decoding parameters. Surprisingly, with proper decoding parameters, the base model alone is already capable of following instructions. See the complete results in App. C.2.
Decoding parameters. When computing the performance of Urial, Lin et al. (2024) use decoding parameters (temperature $=0$, top-$p=1$, repetition penalty $=1.15$) which differ from the default ones in MT-Bench (where temperature depends on the topic of the question, top-$p=1$, repetition penalty $=1$). To clarify the influence of these parameters that were not explicitly discussed in Lin et al. (2024), we compute the performance of base LLM, base LLM with Urial, and fine-tuned LLM when varying decoding configurations. In Fig. 1 we fix the repetition penalty to $1.15$, and vary temperature and top-$p$, for the Mistral-7B-v0.2 models. Surprisingly, we notice that with the decoding parameters of Urial (temperature $=0$, top-$p=1$) even the base model without any in-context example achieves reasonable MT-Bench score (6.61 vs 7.00 of Urial). Moreover, the temperature value appears to have the most influence of the performance of the base model. With Urial, almost all configurations provide very similar results, with the exception of temperature $=1$, top-$p=1$ (possibly because it allows the sampling of generated tokens from the tail of the token distribution, thus producing low-quality text). Finally, the aligned model performs similarly across all decoding schemes, with slightly better results than Urial. These results suggest that (1) the decoding parameters are an overlooked factor contributing to the success of in-context alignment, and (2) fine-tuning adjusts the sampling distribution of language models so that even with high-variance decoding configurations the generated text preserves high quality. We provide additional analysis for no repetition penalty (i.e. $=1$) in Fig. 8 and Llama-3.1-8B in Fig. 9 in App. C.2: these extensive experiments further validate, with different base LLMs and settings, that the decoding scheme is a crucial factor for the instruction-following behavior of base models, and also affects, but to lower degree, in-context alignment. Instead, the fine-tuned models are robust to variations in the decoding parameters. Then, for the rest of the experiments we fix the decoding scheme of Urial.
In-context prompt. Next, we want to disentangle the influence of the various elements of Urial on instruction following performance. In Fig. 2 we track the MT-Bench scores when putting in-context all possible subsets of the three demonstrations of Urial (i.e. from 0 to 3 examples are used). Moreover, for each case we test either adding or not the part of the prompt including the rules to follow. First, we find that neither 1st-turn nor 2nd-turn MT-Bench scores are significantly influenced by the addition of the set of rules, in turn highlighting the importance of the examples. Focusing on 1st-turn score (left plot in Fig. 2), we see that increasing the number of in-context demonstration progressively improves the performance of the base model (Mistral-7B-v0.2 in this case): this observation motivates us to add further high-quality in-context demonstrations, see Sec. 3. Finally, the 2nd-turn has the opposite trend, getting degraded by more (single-round) demonstrations.
Figure 2: Influence of the individual components of Urial. We report 1st and 2nd turn MT-Bench score for every subset of the three demonstrations of Urial with Mistral-7B-v0.2. We test each configuration with (“Rule + Examples”) and without (“Examples-Only”) the Urial set of rules in the in-context prompt. We observe a clear increasing trend in 1st-turn score with more examples, but a decrease in the 2nd-turn performance. The set of rules does not seem to influence the results. The randomness for 0 and 3 examples is caused by the small fluctuations in the score of the GPT-4 judge.
2.3 Importance of question-answer matching
Surprisingly, Min et al. (2022) showed that using demonstrations with random labels does not significantly impair the results of ICL on classification and multiple choice tasks. We conduct a similar study for instruction following, see Table 2. Denoting $\{(X_{i},Y_{i})\}_{i}$ the set of in-context examples, with $X_{i}$ the query and $Y_{i}$ the corresponding ground-truth answer, we test several configurations on two sets of $\{(X_{i},Y_{i})\}_{i}$: the 3 Urial examples and the 3 Urial examples with the 3 examples found by our greedy search from Sec. 3.2 (to check the effect of increasing the number of demonstrations).
First, we do not use any demonstration, i.e., the original base models: with the decoding parameters found in the previous section, this already achieves reasonable results. Surprisingly, providing the questions without the answers (second row in Table 2) degrades the performance, while the opposite, answers only, is effective (0.4-0.7 higher scores than the base model). Next, we permute the answers $Y_{i}$s with a circular shift of one position, so that all correct answers are still contained in the prompt but matched with the wrong question: this significantly degrades the performance, especially with more examples (Urial + Greedy Search), e.g. Llama-3.1-8B attains a score of only 2.13 (note that the minimum score is 1). Also, for each question, we sample a new answer from those provided for other instructions in the same category (in-domain): e.g., a question about coding is paired with an answer from a different coding question, ensuring that, while the content may be incorrect, the style remains appropriate. Although worse than the original one, this configuration achieves decent scores. Finally, we sample answers from instructions belonging to different out-of-domain categories, so that even their style does not match what expected for each question: this leads to the worst performance.
These results show that not all the conclusions from Min et al. (2022) apply to ICL used for instruction following. Most importantly, using answers with correct content and, especially, correct style is crucial for the success of ICL. This property becomes even more evident when increasing the number of examples (to 6 instead of 3), and motivates us to use a highly-curated instruction dataset like SkillMix in experiments in the next sections. Finally, this result suggests that ICL for instruction following works differently compared to other less open-ended tasks.
Table 2: Importance of question-answer matching of demonstrations for in-context alignment. We report the 1st-turn MT-Bench score of Mistral-7B-v0.2 and Llama-3.1-8B when varying the structure of the in-context examples $\{(X_{i},Y_{i})\}_{i}$, where $X_{i}$ is the query and $Y_{i}$ is the corresponding ground-truth answer.
Mistral-7B-v0.2 Llama-3.1-8B Mistral-7B-v0.2 Llama-3.1-8B
no demonstrations $6.52$ $6.25$ $6.52$ $6.25$
$X$ only (no $Y$) $5.90$ $4.48$ $5.53$ $5.10$
$Y$ only (no $X$) $\underline{6.94}$ $\underline{6.83}$ $\underline{7.02}$ $\underline{6.98}$
circular shift of $Y$ $5.04$ $5.69$ $5.59$ $2.13$
in-domain random $Y$ $6.24$ $6.26$ $4.63$ $5.49$
out-of-domain random $Y$ $4.13$ $3.73$ $1.50$ $4.36$
correct $Y$ 6.99 6.95 7.43 7.81
3 Many-Shot In-Context Learning for Instruction Following
Sec. 2.1 indicate that Urial alone is not able, in most cases, to reach the performance of instruct models. In the following, we explore whether we can close this gap. We focus on Mistral-7B-v0.2 (Jiang et al., 2023) and Llama-3.1-8B (Dubey et al., 2024) since (a) both Urial and the aligned model achieve competitive performance in Table 1, and (b) these base models have context windows of 32k and 128k respectively, which can fit many examples (around 50 for Mistral-7B-v0.2 and more than 200 for Llama-3.1-8B). Details on the experimental setup are in App. A.
(a) Mistral-7B-v0.2
(b) Llama-3.1-8B
Figure 3: Scaling the number of demonstrations for alignment with ICL on Mistral-7B-v0.2 and Llama-3.1-8B. We measure the alignment performance of different settings using the MT-Bench score. ICL with more demonstrations quickly saturates and does not bridge the performance gap between the base model and its aligned counterpart. In particular, the ICL alignment performance of 3 random examples from the high-quality SkillMix dataset surpasses that of 3 examples from Urial.
3.1 Scaling up the number of in-context demonstrations
Given the success of many-shot ICL (Agarwal et al., 2024), we test the effect of increasing the number of in-context demonstrations. We sample random examples from the SkillMix dataset, since it contains high-quality question-answer pairs. We consider two scenarios: (1) all demonstrations are randomly chosen from SkillMix (the set of rules from Urial is kept), and (2) we add the new demonstrations on top of the Urial examples. In Fig. 3, we report the results on MT-Bench when varying the number of demonstrations. We repeat sampling with 5 random seeds and show mean and standard deviation over the corresponding results. We observe that for both base LLMs, the 1st-turn score increases with 3, 10 and 20 examples, but then plateaus without clear benefits from scaling up beyond 30 demonstrations (for Llama-3.1-8B the trend becomes even slightly negative with more than 100 examples). Thus, we find that adding up to 30 high-quality examples from SkillMix can improve alignment via ICL. This result contrasts with the findings of Lin et al. (2024), which showed that Urial performed better with 3 demonstrations than with 8.
Next, we observe that in-context alignment does not improve the second-turn score. In fact, adding more examples continues to decrease performance, even falling below that of the base models. This behavior is likely due to the presence of only single-turn examples in the prompt, causing the LLM to respond in the same style without adapting to more complex conversations. Overall, these results show that, unlike in the setting of Agarwal et al. (2024); Bertsch et al. (2024), simply scaling the number of ICL examples is not sufficient to consistently improve the instruction following performance.
Finally, we notice that Llama-3.1-8B using three examples from SkillMix attains, on average, better performance than using the three examples of Urial. Overall, there is no significant difference in the scaling behavior between using or not using the Urial demonstrations. Thus, we conclude that the success of in-context alignment is only partially dependent on the demonstrations themselves, provided they are of sufficient quality (see also Sec. 4).
(a) Search for 4th example
(b) Search for 5th example
(c) Search for 6th example
Figure 4: The distribution of the 1st-turn MT-Bench score (GPT-4-Turbo as judge) on Llama-3.1-8B obtained by adding multiple instructions from SkillMix (Kaur et al., 2024) as a 4th (a), 5th (b), 6th (c) demonstration to Urial. The dashed lines of various colors refer to the 1st-turn MT-Bench score of the obtained searching results. A majority of the 4th examples contribute positively to the model’s instruction-following performance, but the improvement quickly diminishes when running the greedy search for 5th and 6th demonstrations.